# Boltzmann Generators for condensed matter via Riemannian Flow Matching

### v1

@@ 1 Introduction
Continuous normalizing flows (CNFs) (Chen et al., 2018) trained with the scalable and efficient flow-matching training paradigm (Lipman et al., 2023; Albergo and Vanden-Eijnden, 2023; Liu et al., 2023) have emerged as the state-of-the-art approach for modeling complex biomolecular systems (Klein and Noe, 2024; Rehman et al., 2025). However, the development of flexible CNFs for condensed-phase systems, commonly modeled using periodic boundary conditions, remains limited. Previous studies have primarily employed architecturally constrained coupling-flow models Dinh et al. (2017) for crystals (Wirnsberger et al., 2022; Schebek et al., 2024; Ahmad and Cai, 2022) and liquids (Wirnsberger et al., 2020; Coretti et al., 2025), but these approaches were limited by computational cost, restricting training to system sizes of a few hundred particles, which is significantly smaller than what is required to eliminate finite-size effects (Polson et al., 2000).
The goal of this work is to introduce CNFs with periodic boundary conditions in combination with Riemannian flow matching (Chen and Lipman, 2024) and tractable density estimation as a flexible, scalable, and transferable framework for equilibrium sampling of condensed-phase systems. We show that this approach overcomes the scalability limitations of prior coupling-flow architectures by training on a monatomic ice system more than four times larger than previous benchmarks, achieving excellent sample quality while matching the computational budget of the smaller baselines.
@@ 2 Methods
Boltzmann Generators combine exact-likelihood deep generative models, such as normalizing flows (Rezende and Mohamed, 2015; Papamakarios et al., 2021), with importance reweighting to sample from a target Boltzmann distribution $\mu(\mathbf{x})=Z^{-1}\exp[-\beta U(\mathbf{x})]$, which is defined by the potential $U(\mathbf{x})$, $\beta=(k_{B}T)^{-1}$, and the partition function $Z$, which is related to the free energy by $F=-k_{B}T\log Z$. The model is trained to transform a prior distribution $p(\mathbf{x})$ into a proposal distribution $\tilde{p}(\mathbf{x})$ that approximates $\mu(\mathbf{x})$, via an invertible mapping $f_{\theta}$. Unbiased estimates can be obtained by reweighting the generated samples to the target distribution using the importance weights
$w(\mathbf{x})=\exp\left(\beta_{p}U_{p}(\mathbf{x})-\beta_{\mu}U_{\mu}(f_{\theta}(\mathbf{x}))+\log\left|\det J_{f}(\mathbf{x})\right|\right),$ (1)
@@
where $J_{f}$ denotes the Jacobian of the transformation. The importance weights further provide a measure of modeling performance via the effective sample size (ESS) defined by ${\rm ESS}=(\sum_{i=1}^{N_{S}}w(\mathbf{x}^{(i)}))^{2}/\sum_{i=1}^{N_{S}}(w(\mathbf{x}^{(i)}))^{2}$. Within the framework of targeted free energy perturbation (TFEP) (Jarzynski, 2002), a trained BG can be used to directly estimate the free energy difference between prior and target states through $\beta\Delta F_{p\mu}=-\log\mathbb{E}_{\mathbf{x}\sim p(\mathbf{x})}\left[w(\mathbf{x})\right]$, without needing to sample intermediate states via MD simulations. Knowing the free energy of the prior state allows estimation of the absolute free energy of the target system and, consequently, a comparison of the thermodynamic stability of different system states.
@@
Normalizing Flows learn an invertible mapping $f_{\theta}$ to transform samples from a prior distribution $p(\mathbf{x})$ to a generated density $\tilde{p}(\mathbf{x})$ that approximates the target distribution $\mu(\mathbf{x})$. CNFs (Chen et al., 2018) model this transformation in continuous time by solving an ordinary differential equation (ODE) over $t\in[0,1]$,
$\frac{df_{\theta}^{t}(\mathbf{x})}{dt}=v_{\theta}(t,f_{\theta}^{t}(\mathbf{x})),\quad f_{\theta}^{0}(\mathbf{x})=\mathbf{x},$ (2)
@@
where $v_{\theta}$ is a time-dependent vector field with learnable parameters $\theta$. The resulting density evolution follows the continuous change-of-variables formula, where the log-determinant of the Jacobian corresponds to the integrated divergence of the vector field. CNFs can be naturally extended to general Riemannian manifolds $\mathcal{M}$ by formulating and solving the associated ODE directly on $\mathcal{M}$.
@@
Riemannian Flow Matching. Lipman et al. (2023) introduced flow matching for efficient and simulation-free training of CNFs, which has been generalized to manifolds by Chen and Lipman (2024). In the case of periodic boundary conditions, corresponding to the flat torus, the Riemannian conditional flow matching (RFM) objective allows to train the vector field through
@@
where $p_{t}(\mathbf{x}|\mathbf{z})$ is a probability path conditioned on an arbitrary distribution $\mathbf{z}$. A proven choice for this path and the conditional distribution $\mathbf{z}$ is the direct constant speed interpolation between a single prior and target sample, which yields the conditional vector field
@@ 3 Implementation
Symmetries and Prior Alignment. Embedding symmetries of the equilibrium density, including invariance to global translations, rotations, and particle permutations, into generative flow models has been demonstrated to substantially improve data efficiency and sample quality (Köhler et al., 2020; Satorras et al., 2021; Klein et al., 2023). In principle, the presented framework allows for straightforward incorporation of such symmetries, by combining a symmetry-invariant prior with an equivariant vector field. However, for the specific case of crystalline solids, we exploit the underlying lattice structure to adopt a distinct, physics-informed strategy: A 3N-dimensional wrapped Gaussian centered at equilibrium lattice sites (Frenkel and Ladd, 1984; Wirnsberger et al., 2022) (see Appendix A.2.1).
As the lattice prior is not translation invariant, we restrict the system to a mean-free subspace, by fixing the center of mass of both the prior and target data. The learned dynamics are constrained to preserve the center of mass by subtracting the mean particle velocity from both the learned vector field and the conditional target field. The continuous $\mathrm{SO}(3)$ symmetry of Euclidean systems is reduced to a discrete subgroup for periodic boundary conditions. Our lattice-based prior implicitly enforces a global alignment with the simulation box axes and a trivial assignment between prior and target samples. While we perform mini-batch optimal transport reordering (Tong et al., 2023; Pooladian et al., 2023), we omit further searches over discrete rotations and only adapt the transport cost calculation to torus distances.
Vector Field Parameterization. To scale effectively to larger systems, we adopt a local modeling approach for the vector field (Schebek et al., 2025b). We leverage the architectural flexibility of CNFs to adapt a powerful, size-transferable representation of atomic environments commonly used in machine learning interatomic potential architectures — specifically the Equivariant Transformer (ET) (Pelaez et al., 2024) as proposed in Hassan et al. (2024) — and consequently term our method RFM-ET in the following. Details on adaptations and hyperparameters are in Appendix A.5.
@@
As the chosen vector field parametrization naturally incorporates the toroidal topology and its outputs are directly in the Euclidean tangent space $\mathcal{T}_{\mathbf{x}}\mathcal{M}=\mathbb{R}^{3N}$, the Riemannian divergence simplifies to the standard Euclidean trace of the Jacobian, and we can use off-the-shelf ODE solvers, simply by projecting the final state back onto the torus. The effective $\mathcal{O}(N^{2})$ complexity of calculating the exact divergence renders it infeasible for large systems. Hutchinson’s trace estimator (Hutchinson, 1989) offers an $\mathcal{O}(N)$ alternative but is often considered unsuitable for Boltzmann sampling applications due to noise and statistical bias. Although Hutchinson’s method provides an unbiased estimate of the density change in a CNF, thermodynamic reweighting relies on the exponentiated change of that density. Due to Jensen’s inequality (Jensen, 1906), the nonlinearity of the exponential function transforms the estimator’s variance into a systematic bias in the importance weights. This bias propagates to all downstream equilibrium estimates, such as free energy differences. To address this, we propose a bias correction based on the second-order cumulant expansion (Zwanzig, 1954). We estimate the variance of the integrated divergence, $\hat{\sigma}^{2}$, and apply a correction term of $\frac{1}{2}\hat{\sigma}^{2}$ directly to the log-weights. This effectively removes the bias introduced by the estimator noise, allowing for accurate thermodynamic estimates. A more detailed discussion is provided in Appendix A.3.
@@ 4 Experimental Results
Figure 2: Left: Effective sample size computed for the models described in the text, including results for models evaluated on the same system size as used during training, as well as for models transferred to larger system sizes than those seen during training. Error bars are shown where visible. Right: The deviation to the reference free energy of a 512 particles mW system in the cubic ice phase, computed with a varying number of Hutchinson probes, with and without bias correction. The shaded area corresponds to the accuracy needed for precise phase separation between hexagonal and cubic ice. Errors for our method in all plots were estimated over 1010k samples obtained from three models trained with different seeds.
@@
We evaluate the performance of our method with respect to both the quality of the generated samples and its accuracy as a free energy estimator. As a test system, we consider the common benchmark system cubic ice in the monoatomic water (mW) (Molinero and Moore, 2009) parametrization. Figure 1 (left and center) shows radial distribution functions and histograms of the potential energy as obtained from MD and RFM-ET for system sizes of 216 and 1000 particles. Importantly, even without any reweighting, excellent agreement is achieved for both observables, demonstrating that RFM-ET can accurately approximate the target Boltzmann distribution.
@@
The strong performance of RFM-ET is further reflected in the high ESS values shown in the left panel of Fig. 2. These values, obtained from models evaluated on the same system size they were trained on, are compared to those of a global transformer-based coupling flow (GCF) from Wirnsberger et al. (2022), as well as the joint (LCF-J) and marginal (LCF-M) distributions of a local augmented coupling flow Schebek et al. (2025b). RFM-ET consistently matches or outperforms the others, demonstrating its high sampling quality. Notably, training on a system size of $N=512$ yields a sampling efficiency two orders of magnitude higher than that of GCF, and, unlike the other models for which $N=1000$ is computationally infeasible, RFM-ET still achieves an ESS above three percent on this system size. Importantly, RFM-ET significantly reduces training time and enables training on large system sizes, such as $N=1000$, using the same computational budget that LCF requires to train on only $N=216$ (Fig. 1, right).
@@
A crucial feature of our method is its transferability to larger system sizes, allowing models trained on small systems to be evaluated on larger systems without requiring additional data, as shown in the center of Fig. 2. While RFM-ET matches the sampling efficiency of LCF-M when transferring models trained on $N=216$ to $N=512$ and $N=1000$, only RFM-ET can be trained on $N=512$, which allows increasing the sampling efficiency for $N=1000$ by more than fivefold.
@@
Figure 2 (right) demonstrates the efficiency of the proposed bias correction by showing the relationship between the number of Hutchinson probes used per integration step and the free energy estimate obtained via TFEP. While the variance of the free energy estimates decreases rapidly with increasing number of probes, systematic bias remains the dominant source of error when no correction is applied. Even with 256 probes, the free energy estimated using biased weights is underestimated. In contrast, the bias-corrected weights yield a highly accurate estimate with as few as 16 probes.
@@ 5 Discussion
We showed that Riemannian Flow matching provides a powerful tool for equilibrium sampling of condensed matter, exhibiting superior scaling during training compared to coupling flows, enabling training on system sizes that were previously infeasible. Our method not only improves sampling on directly trained systems but also enhances performance when models are transferred to larger systems. Two primary challenges remain. First, inference latency is dominated by the ODE integration and the stochastic divergence estimation required for density evaluation. Generating samples, and particularly computing their corresponding weights, is substantially faster with coupling flows: generating 10,000 samples for the 216-particle system with 10 integration steps and one Hutchinson probe takes roughly 10 minutes on a single A5000, with time scaling linearly with both the number of probes and system size. Second, although flow-matching formulations are generally robust even with limited training data (Schebek et al., 2025a), the method still relies on high-quality configurations of the target distribution, even though data-free training of continuous-time models is being actively explored (Havens et al., 2025).

### v2

@@ 1 Introduction
Continuous normalizing flows (CNFs) (Chen et al., 2018) trained with the scalable and efficient flow-matching training paradigm (Lipman et al., 2023; Albergo and Vanden-Eijnden, 2023; Liu et al., 2023) have emerged as the state-of-the-art approach for modeling complex biomolecular systems (Klein and Noe, 2024; Rehman et al., 2025). However, the development of flexible CNFs for condensed-phase systems, commonly modeled using periodic boundary conditions, remains limited (Grenioux et al., 2025). Previous studies have primarily employed architecturally constrained coupling-flow models (Dinh et al., 2017) for crystals (Wirnsberger et al., 2022; Schebek et al., 2024; Ahmad and Cai, 2022) and liquids (Wirnsberger et al., 2020; Coretti et al., 2025), but these approaches were limited by computational cost, restricting training to system sizes of a few hundred particles, which is significantly smaller than what is required to eliminate finite-size effects (Polson et al., 2000).
The goal of this work is to introduce CNFs with periodic boundary conditions in combination with Riemannian flow matching (Chen and Lipman, 2024) and tractable density estimation as a flexible, scalable, and size-transferable framework for equilibrium sampling of condensed-phase systems. We show that this approach overcomes the scalability limitations of prior coupling-flow architectures by training on a monatomic ice system more than four times larger than previous benchmarks, achieving excellent sample quality while matching the computational budget of the smaller baselines.
@@ 2 Methods
Boltzmann Generators combine exact-likelihood deep generative models, such as normalizing flows (Rezende and Mohamed, 2015; Papamakarios et al., 2021), with importance reweighting to sample from a target Boltzmann distribution $\mu(\mathbf{x})=Z^{-1}\exp[-\beta U(\mathbf{x})]$. This distribution is defined by the potential $U(\mathbf{x})$, the inverse temperature $\beta=(k_{B}T)^{-1}$, with Boltzmann constant $k_{B}$, temperature $T$ and the partition function $Z$. The model is trained to learn an invertible mapping $f_{\theta}$ that transforms a prior sample $\mathbf{x}_{0}\sim p(\mathbf{x}_{0})$ into a generated sample $\mathbf{x}_{1}=f_{\theta}(\mathbf{x}_{0})$, whose distribution $\tilde{p}(\mathbf{x}_{1})$ approximates $\mu(\mathbf{x}_{1})$. Unbiased estimates can be obtained by reweighting the generated samples to the target distribution using the importance weights
$w(\mathbf{x}_{0})=\exp\left(\beta_{p}U_{p}(\mathbf{x}_{0})-\beta_{\mu}U_{\mu}(f_{\theta}(\mathbf{x}_{0}))-\Delta\log p\right),$ (1)
such that $w(\mathbf{x}_{0})\propto\mu(f_{\theta}(\mathbf{x}_{0}))/\tilde{p}(f_{\theta}(\mathbf{x}_{0}))$. The density change $\Delta\log p$ is defined as
$\Delta\log p:=\log\tilde{p}(f_{\theta}(\mathbf{x}_{0}))-\log p(\mathbf{x}_{0})=-\log\left|\det J_{f}(\mathbf{x}_{0})\right|,$ (2)
@@
where $J_{f}$ denotes the Jacobian of the transformation. The importance weights further provide a measure of modeling performance via the effective sample size (ESS) (Kish, 1965) defined by ${\rm ESS}=(\sum_{i=1}^{N_{S}}w(\mathbf{x}_{0}^{(i)}))^{2}/\sum_{i=1}^{N_{S}}(w(\mathbf{x}_{0}^{(i)}))^{2}$, which is typically reported relative to the number of samples $N_{S}$. While the free energy $F$, which governs phase stability, is defined in terms of the partition function $Z$ as $F=-k_{B}T\log Z$, evaluating the partition function is typically intractable. Within the framework of targeted free energy perturbation (TFEP) (Jarzynski, 2002), a trained BG can be used to directly estimate the free energy difference between prior and target states through $\beta\Delta F_{p\mu}=-\log\mathbb{E}_{\mathbf{x}_{0}\sim p(\mathbf{x}_{0})}\left[w(\mathbf{x}_{0})\right]$. This allows the calculation of free-energy differences without sampling intermediate states via MD simulations, which are required by traditional estimators that rely on sufficient phase-space overlap between distributions (Shirts and Chodera, 2008). Knowing the free energy of the prior state allows estimation of the absolute free energy of the target system and, consequently, a comparison of the thermodynamic stability of different phases.
@@
Normalizing Flows learn an invertible mapping $f_{\theta}$ to transform samples from a prior distribution $p(\mathbf{x}_{0})$ to a generated density $\tilde{p}(\mathbf{x}_{1})$ that approximates the target distribution $\mu(\mathbf{x}_{1})$. CNFs (Chen et al., 2018) model this transformation in continuous time by solving an ordinary differential equation (ODE) over $t\in[0,1]$,
$\frac{df_{\theta}^{t}(\mathbf{x})}{dt}=v_{\theta}(t,f_{\theta}^{t}(\mathbf{x})),\quad f_{\theta}^{0}(\mathbf{x})=\mathbf{x}_{0},$ (3)
@@
where $v_{\theta}$ is a time-dependent vector field with learnable parameters $\theta$. The resulting density evolution follows the continuous change-of-variables formula
$\Delta\log p=-\int_{0}^{1}\nabla\cdot v_{\theta}(t,f_{\theta}^{t}(\mathbf{x}_{0}))\,dt,$ (4)
where the divergence of the vector field is equivalent to the trace of the vector field’s Jacobian. CNFs can be naturally extended to general Riemannian manifolds $\mathcal{M}$ by formulating and solving the associated ODE directly on $\mathcal{M}$ (Mathieu and Nickel, 2020).
@@
Riemannian Flow Matching. While normalizing flows are commonly trained by maximum-likelihood objective, Lipman et al. (2023) introduced flow matching for efficient and simulation-free training of CNFs, which has been generalized to manifolds by Chen and Lipman (2024). In the case of periodic boundary conditions, corresponding to the flat torus, the Riemannian conditional flow matching (RFM) objective allows to train the vector field through
@@
where $p_{t}(\mathbf{x}|\mathbf{z})$ is a probability path conditioned on an arbitrary distribution $\mathbf{z}$. A proven choice for this path and the conditional distribution $\mathbf{z}$ is the direct constant speed interpolation between a single prior and target sample (Chen and Lipman, 2024), which yields the conditional vector field
@@ 3 Implementation
Symmetries and Prior Alignment. The equilibrium configurational distribution naturally inherits the symmetries of the potential energy, including invariance wrt. global translations, rotations, and particle permutations. Embedding those symmetries into generative flow models has been demonstrated to substantially improve data efficiency and sample quality (Köhler et al., 2020; Satorras et al., 2021; Klein et al., 2023). In principle, the presented framework allows for straightforward incorporation of such symmetries, by combining a symmetry-invariant prior, such as the uniform distribution on the torus, with an equivariant vector field, as done in Grenioux et al. (2025). However, for the specific case of crystalline solids, we exploit the underlying lattice structure to adopt a distinct, physics-informed strategy: A 3NN-dimensional wrapped Gaussian centered at equilibrium lattice sites (Wirnsberger et al., 2022) (see Appendix A.2.1) known as Einstein crystal (Frenkel and Ladd, 1984).
As the chosen lattice prior is not translation invariant, we restrict the system to a mean-free subspace, by fixing the center of mass of both the prior and target data. The learned dynamics are constrained to preserve the center of mass by subtracting the mean particle velocity from both the learned vector field and the conditional target field. The continuous $\mathrm{SO}(3)$ symmetry of Euclidean systems is reduced to a discrete subgroup for periodic boundary conditions. Our lattice-based prior implicitly enforces a global alignment with the simulation box axes and a trivial assignment between prior and target samples. While we perform mini-batch optimal transport reordering (Tong et al., 2023; Pooladian et al., 2023), we omit further searches over discrete rotations and only adapt the transport cost calculation to torus distances.
Vector Field Parameterization. To scale effectively to larger systems, we adopt a local modeling approach for the vector field, which ensures that the learned dynamics are transferable across different system sizes (Schebek et al., 2025b). We leverage the architectural flexibility of CNFs to adapt a powerful, size-transferable representation of atomic environments commonly used in machine learning interatomic potential architectures — specifically the Equivariant Transformer (ET) (Pelaez et al., 2024) as proposed in Hassan et al. (2024) — and consequently term our method RFM-ET in the following. Details on adaptations and hyperparameters are in Appendix A.5.
@@
As the chosen vector field parametrization naturally incorporates the toroidal topology and its outputs are directly in the Euclidean tangent space $\mathcal{T}_{\mathbf{x}}\mathcal{M}=\mathbb{R}^{3N}$, the Riemannian divergence simplifies to the standard Euclidean trace of the Jacobian, and we can use off-the-shelf ODE solvers, simply by projecting the final state back onto the torus. Due to the GNN’s local cutoff, the computational cost of evaluating the CNF vector field scales linearly with system size, $\mathcal{O}(N)$. However, calculating the exact divergence requires computing the trace of the full Jacobian. Because the input dimension $D$ is proportional to $N$, obtaining this trace requires $\mathcal{O}(N)$ evaluations of the vector field. This results in an overall $\mathcal{O}(N^{2})$ complexity for both computation and memory, making the exact calculation infeasible for large systems.
Hutchinson’s trace estimator (Hutchinson, 1989) offers an $\mathcal{O}(N)$ alternative but is often considered unsuitable for Boltzmann sampling applications due to noise and statistical bias. Although Hutchinson’s method provides an unbiased estimate of the vector field’s divergence and thus the log-density change $\Delta\log p$, thermodynamic reweighting relies on the exponentiated log-density change. Due to Jensen’s inequality (Jensen, 1906), the nonlinearity of the exponential function transforms the estimator’s variance into a systematic bias in the importance weights. This bias propagates to all downstream equilibrium estimates, such as free energy differences. To address this, we propose a bias correction based on the second-order cumulant expansion (Zwanzig, 1954). By treating the stochastic error from the trace estimator as a fluctuating work term, we can expand the expectation of the likelihood estimate. Because the total accumulated noise arises from the sum of independent stochastic evaluations over multiple integration steps, its distribution rapidly converges to a Gaussian via the central limit theorem. Consequently, cumulants higher than the second order vanish, justifying the truncation of the series. For each sample, we estimate the variance of the stochastic log-density change, $\hat{\sigma}^{2}$, and apply a correction term directly to the log-weights:
$\widehat{\log w_{\text{corr}}}(\mathbf{x}_{0})\coloneq\widehat{\log w}(\mathbf{x}_{0})-\frac{1}{2}\hat{\sigma}^{2}(\mathbf{x}_{0}).$ (7)
This effectively removes the bias introduced by the estimator noise, allowing for accurate thermodynamic estimates. To efficiently evaluate the correction term, we augment the ODE state with the empirical variance of the trace estimator, integrating it alongside the positions and log density. A more detailed discussion is provided in Appendix A.3 and Appendix A.4.
@@ 4 Experimental Results
Figure 2: Left: Effective sample size computed for the models described in the text, including results for models evaluated on the same system size as used during training, as well as for models transferred to larger system sizes than those seen during training. Error bars are shown where visible. Right: The deviation to the reference free energy of a 512 particles mW system in the cubic ice phase, computed with a varying number of Hutchinson probes, with and without bias correction. The gray shaded area indicates a tolerance window of $10^{-4}$. Errors for our method in all plots were estimated over 1010k samples obtained from three models trained with different seeds.
@@
We evaluate the performance of our method regarding both the quality of the generated samples and its accuracy as a free energy estimator. As a test system, we consider the common benchmark system cubic ice in the monoatomic water (mW) (Molinero and Moore, 2009; Wirnsberger et al., 2022) parametrization. Figure 1 (left and center) shows radial distribution functions and histograms of the potential energy as obtained from MD and RFM-ET for system sizes of 216 and 1000 particles. Importantly, even without any reweighting, excellent agreement is achieved for both observables, demonstrating that RFM-ET can accurately approximate the target Boltzmann distribution.
@@
The strong performance of RFM-ET is further reflected in the high ESS values shown in the left panel of Fig. 2. We first evaluate the performance of our model in the non-transferable setting, with all models assessed on the system size they were trained on. Where available, we compare to literature results obtained using a global transformer-based coupling flow (GCF) (Wirnsberger et al., 2022) and a local augmented coupling flow (LCF) (Schebek et al., 2025b); for the LCF, estimates based on joint (LCF-J) and marginal (LCF-M) densities are included. RFM-ET consistently matches or outperforms the others, demonstrating its high sampling quality. Notably, training on a system size of $N=512$ yields a sampling efficiency two orders of magnitude higher than that of GCF, and, unlike the other models for which $N=1000$ is computationally infeasible, RFM-ET still achieves an ESS above three percent on this system size. Importantly, RFM-ET significantly reduces training times and enables training on large system sizes, such as $N=1000$, using the same computational budget that LCF requires to train on only $N=216$ (Fig. 1, right).
@@
A crucial feature of our method is its transferability to larger system sizes, allowing models trained on small systems to be evaluated on larger systems without requiring additional data or training, similarly to the LCF architecture. As shown in the center of Fig. 2, RFM-ET matches the sampling efficiency of LCF-M when transferring models trained on $N=216$ to $N=512$ and $N=1000$, but only RFM-ET can be trained efficiently on $N=512$, which allows increasing the sampling efficiency for $N=1000$ by more than fivefold.
@@
The right panel of Fig. 2 demonstrates the efficiency of the proposed bias correction by showing the relationship between the number of Hutchinson probes used per integration step and the free energy estimate obtained via TFEP. While the variance of the free energy estimates decreases rapidly with increasing number of probes, systematic bias remains the dominant source of error when no correction is applied. Even with 256 probes, the free energy estimated using biased weights is underestimated. In contrast, with as few as 16 probes, the bias-corrected weights allow the free energy estimates to converge to an accuracy of around $10^{-4}$, enabling resolution of free energy differences on the order of $10^{-3}$ between hexagonal and cubic ice under the given settings (Schebek et al., 2025b). With fewer probes, the variance of the log-density change is underestimated, resulting in an insufficient bias correction.
@@ 5 Discussion
We showed that Riemannian Flow matching provides a powerful tool for equilibrium sampling of condensed matter, exhibiting superior scaling during training compared to coupling flows, enabling training on system sizes that were previously infeasible. Our method improves sampling on directly trained systems and enhances performance when models are transferred to larger systems. While the results presented here focus on cubic ice, we recently showed that the methodology is also applicable to other crystal structures within the mW potential and to systems described by the Lennard-Jones potential (Schebek et al., 2025a). As also noted in the same study, two primary challenges remain. First, inference cost is high due to the computationally expensive ODE integration and stochastic divergence estimation required for density evaluation, making it one to two orders of magnitude slower than sampling with coupling flows. Second, although flow-matching formulations are generally robust even with limited training data, the method still relies on high-quality configurations of the target distribution, preventing the amortization of training costs through generalization over external conditions (Schebek et al., 2024). A promising direction is data-free training of continuous-time models, which is currently being actively explored (Havens et al., 2025; Blessing et al., 2026). Future work could further explore the integration of consistency distillation (Geng et al., 2025) to enable few-step sampling without sacrificing the benefits of the continuous-time formulation, as well as the extension to the $NPT$ ensemble and physically more realistic potentials.